OpenStudy (anonymous):

If tan(θ) = − 8/ 9 and sin(θ) > 0, find the exact values of each of the trig functions how would you work this out? could someone please explain?

OpenStudy (anonymous):

\[ \tan ^2(\theta )+1=\left(\frac{8}{9}\right )^2+1=\sec ^2(\theta ) \] Solve and you find \[ \sec(\theta)=\pm\frac{\sqrt{145}}{9} \] Which sign do you choose?

OpenStudy (anonymous):

You choose the negative sign since \( \cos(\theta)<0\)

OpenStudy (anonymous):

\[\Large \cos(\theta)=-\frac{9}{\sqrt {145}} \]

OpenStudy (anonymous):

Can you find the remaining?

OpenStudy (anonymous):

i dont know how to find them. havent done these questions before

OpenStudy (anonymous):

Solve this to find \(\sin(\theta)\) \[ \tan (\theta )=\frac{\sin (\theta )}{\cos (\theta )}=-\frac{8}{9}=-\frac{\sin (\theta )}{\frac{9}{\sqrt{145}}} \\ \Large \sin (\theta )=\frac{8}{\sqrt{145}}\\ \Large \csc(\theta)=\frac{\sqrt{145} }{8} \]

OpenStudy (anonymous):

\[ \Large \cot(\theta)=-\frac 98 \]